Suppose that 55% of all adults regularly consume coffee, 45% regularly consume carbonated soda, and 70% regu- larly consume at least one of these two products. a. What is the probability that a randomly selected 14. adult regularly consumes both coffee and soda? b. What is the probability that a randomly selected adult doesnt regularly consume at least one of these two products?
Answer #1

Solution:-

14)

a) The probability that a randomly selected adult regularly consumes both coffee and soda is 0.30.

P(Coffee) = 0.55

P(Soda) = 0.45

P(Coffee or Soda) = 0.70

P(Coffee or Soda) = P(Coffee) + P(Soda) - P(Coffee and Soda)

0.70 = 0.55 + 0.45 - P(Coffee and Soda)

P(Coffee and Soda) = 1.0 - 0.70

P(Coffee and Soda) = 0.30

b) The probability that a randomly selected adult doesn't regularly consumes atleast one of these two products is 0.30.

P(Not Coffee or Soda) = 1 - P(Coffee or Soda)

P(Not Coffee or Soda) = 1 - 0.70

P(Not Coffee or Soda) = 0.30